OpenStudy (anonymous):

Someone examined 300 bulbs found 22 defective.At this rate,how many defective bulbs would there be in 33,000 bulbs?

jimthompson5910 (jim_thompson5910):

solve for x 22/300 = x/33000 22*33000 = 300x ... ... ... x = ???